# Dynamic Sparse Training versus Dense Training: The Unexpected Winner in Image Corruption Robustness

### v1

@@ Abstract
It is generally perceived that Dynamic Sparse Training opens the door to a new era of scalability and efficiency for artificial neural networks at, perhaps, some costs in accuracy performance for the classification task. At the same time, Dense Training is widely accepted as being the ”de facto” approach to train artificial neural networks if one would like to maximize their robustness against image corruption. In this paper, we question this general practice. Consequently, we claim that, contrary to what is commonly thought, the Dynamic Sparse Training methods can consistently outperform Dense Training in terms of robustness accuracy, particularly if the efficiency aspect is not considered as a main objective (i.e., sparsity levels between 10% and up to 50%), without adding (or even reducing) resource cost. We validate our claim on two types of data, images and videos, using several traditional and modern deep learning architectures for computer vision and three widely studied Dynamic Sparse Training algorithms. Our findings reveal a new yet-unknown benefit of Dynamic Sparse Training and open new possibilities in improving deep learning robustness beyond the current state of the art.
@@ 1 Introduction
If one would like to maximize as much as possible the robustness of a deep learning model on noisy more realistic data (e.g., image corruption, out of distribution), what would be the typical approach? Would one train directly a dense neural network (referred further as Dense Training) or a sparse neural network (referred further as Sparse Training)? Of course, there are many techniques to enhance the model robustness, but still they would be applied on top of a basic optimization process which usually requires training an artificial neural network. In order to give an answer to the previously posed question, we would need to quantify the research and practitioner community perception, which is a hard problem in itself. Yet, given the large body of literature on the topic, it is safely to assume that the usual answer would be ”Yes, Dense Training is the typical solution to maximize the model robustness.”. Nevertheless, the reader can answer to this question by themselves, while further in this manuscript we would like to bring more clarity and some new perspectives on this topic.
@@
By connecting the dots, an intuitive research question arises: Is somehow DST implicitly improving model robustness, and no one knew about this yet?
Although several studies have investigated the impact of pruning model from Dense Training, lottery ticket hypothesis methods Frankle & Carbin (2019); Frankle et al. (2020); Renda et al. (2020) on image corruption robustness Liebenwein et al. (2021); Diffenderfer et al. (2021) and adversarial robustness Liao et al. (2022), other research shows that static sparsified networks, when increased width and depth to maintain capacity, can also improve robustness Timpl et al. (2022). However, a closer look at current DST research reveals a literature gap: the overlooked evaluation of DST models under various unexpected image corruptions. In an attempt to respond to this, we examined one of the classical DST algorithms, Sparse Evolutionary Training (SET) Mocanu et al. (2018), as illustrated in Figure 1 (Left). We find that SET (with a sparsity ratio of $0.5$) can markedly improve model robustness against 18 out of 19 image corruptions, compared to Dense Training. In light of the strong performance of the DST method (e.g., SET) against common corruptions , we propose the Dynamic Sparsity Corruption Robustness (DSCR) Hypothesis.
@@ 2.2 Dynamic Sparse Training
Dynamic Sparse Training (DST) Mocanu et al. (2018); Bellec et al. (2018); Mostafa & Wang (2019); Evci et al. (2020); Chen et al. (2021); Yuan et al. (2021); Liu et al. (2021a), as a sparse-to-sparse training paradigm, starts from scratch and maintains partial parameters throughout training, providing a promising approach to significantly improve training and inference efficiency while still preserving accuracy. Recent research has focused on deciphering the mechanisms of DST Liu et al. (2021c); Evci et al. (2022; 2019), and has broadened its application in various domains, such as continuous learning Sokar et al. (2023), reinforcement learning Graesser et al. (2022); Tan et al. (2023), feature selection Sokar et al. (2022); Atashgahi et al. (2022), time series classification Xiao et al. (2022), network architecture design Liu et al. (2023), and vision transformers Chen et al. (2021).
Moreover, recent studies have concentrated on developing improved DST algorithms Dettmers & Zettlemoyer (2019); Mostafa & Wang (2019); Yuan et al. (2021); Jayakumar et al. (2020); Liu et al. (2021b). For instance, Dettmers & Zettlemoyer (2019) introduced a technique to regrow weights based on the momentum of existing weights, while Yuan et al. (2021) proposed a method that utilizes both weight and gradient information for pruning. More recently, it has been demonstrated that DST can enhance the performance of DNNs on biased data Zhao et al. (2023) and adversarial samples Chen et al. (2022b). However, the majority of DST research relies on pristine benchmark datasets, leaving the robustness of DST models against image corruption relatively unexplored. This paper aims to systematically investigate the robustness behavior of DST when exposed to various image corruptions.
@@ 3.1 Problem Definition
Given the i.i.d. training samples $\left\{\left(\bm{x}_{i}^{\operatorname{tr}},y_{i}^{\operatorname{tr}}\right)\right\}_{i=1}^{n_{\text{tr }}}$ from a training distribution $p_{\operatorname{tr}}(\bm{x},y)$, our goal is to predict the labels $\hat{y}_{i}^{\text{te}}$ of test examples $\left\{\left(\bm{x}_{i}^{\text{te }},y_{i}^{\text{te}}\right)\right\}_{i=1}^{n_{\text{te}}}$ drawn from a test distribution $p_{\text{te }}(\bm{x},y)$. We decompose $p_{\operatorname{tr}}(\bm{x},y)$ and $p_{\text{te }}(\bm{x},y)$ into the marginal distribution and the conditional probability distribution, that is:
$p_{\operatorname{tr}}(\bm{x},y)=p_{\operatorname{tr}}(\bm{x})p_{\operatorname{tr}}(y\mid\bm{x}),$
$p_{\text{te }}(\bm{x},y)=p_{\text{te }}(\bm{x})p_{\text{te}}(y\mid\bm{x}).$
$p_{\text{tr }}(\bm{x})\neq p_{\text{te }}(\bm{x})$, and $\left.p_{\operatorname{tr}}(y\mid\bm{x})=p_{\mathrm{te}}(y\mid\bm{x}\right)$, in our setting. To evaluate the performance of the models in the aforementioned problem setting, we utilize standard image benchmark datasets such as CIFAR10 Krizhevsky et al. (2009), CIFAR100 Krizhevsky et al. (2009), TinyImageNet Le & Yang (2015) and ImageNet Deng et al. (2009) for training. Then, we evaluate the performance (robustness accuracy) of the models on the corresponding corruption benchmark datasets, including CIFAR10-C, CIFAR100-C, TinyImageNet-C and ImageNet- C/$\bar{\mathbf{C}}$/3DCC. More details of datasets can be found in Appendix A.2.
@@ 3.2 Dynamic Sparse Training
  • DST with soft memory budget in forward: Memory-Economic Sparse Training (MEST) Yuan et al. (2021), with Soft Memory Bound Elastic Mutation, starts with a sparse model having a memory budget of $\frac{\|\bm{\theta}^{\prime}\|_{0}}{\|\bm{\theta}\|_{0}}=b+b_{s}$, where $b_{s}>0$ represents the soft memory bound that decays during training. Subsequently, at intervals of $\Delta t$, weights are removed based on magnitude and gradient information, reducing the memory budget to $b$, and then regrown it to $b+b_{s}$, as illustrated in Figure 1 (right (c)). The soft memory constraint $b_{s}$, gradually decays to zero, allowing the model to reach the target memory budget $b$ by the end of training.
@@ 4.1 Overall Performance
Figure 2 showcases the average robust accuracy across various types of corruption, with severity levels ranging from 1 to 5, on the CIFAR10/100-C and TinyImageNet-C datasets. We can find that DST models outperform their Dense training counterparts (as indicated by the blue line in Figure 2) in terms of robustness accuracy across multiple sparsity ratios. This trend is evident in both the random regrow (Figure 2, left) and gradient-based regrow (Figure 2, right) strategies. Specifically, for CIFAR100-C and TinyImageNet- C, DST models with sparsity ratios ranging from $0.3$ to $0.7$ generally show encouraging improvements in robustness. For CIFAR10-C, at certain sparsity ratios, such as $0.4$, the overall DST methods achieve decent robust performance against image corruptions. Moreover, based on our experimental setup, at least $40\%$ of both training computational and memory costs can be saved while simultaneously enhancing corruption robustness. Details on FLOPs and parameter counts are available in the Appendix A.3.2. This finding substantiates our DSCR Hypothesis across different DST methods, varying sparsity ratios, and architectures (i.e., VGG, ResNet, and EfficientNet).
@@ 4.2 Performance on Corruption Types
We further evaluate the DST methods across various corruption types with different severities. Figure 3 provides detailed relative robustness gains  [^2] Defined as $({\tt Acc_{s,sl}-\tt Acc_{d,sl}})/{\tt Acc_{d,sl}}$, where $\tt Acc_{s,sl}$ and $\tt Acc_{d,sl}$ represent the accuracies of models trained using DST and dense training, respectively, at a severity level of $\tt sl$. for each type of corruption in ImageNet-C and ImageNet-$\bar{\mathbf{C}}$ Mintun et al. (2021), compared to their dense model counterparts. It is worth mentioning that the Y-axis in Figure 3 represents different types of corruption, which are ordered according to the amount of high-frequency information they contain, as in Saikia et al. (2021). Intriguingly, on ImageNet-C and ImageNet-$\bar{\mathbf{C}}$  [^3] Our experiments primarily focus on DST with a gradient-based regrow strategy when training on ImageNet. As our implementation is based on PyTorch, the gradient-based regrow approach tends to achieve strong performance on ImageNet more quickly with this framework Liu et al. (2021a); Nowak et al. (2023). However, it is worth noting that the random-based approach has also shown significant improvements when implemented in JAX Lee et al. (2023), which we plan to explore in future work., DST models show significant potential in enhancing robustness against specific corruptions, such as impulse, Gaussian, and shot noise on ImageNet-C, as well as brownish, Perlin, plasma, and blue noise on ImageNet-$\bar{\text{C}}$. It is worth noting that these corruptions are typically characterized by relatively high-frequency information.
This finding is consistent on CIFAR10-C, CIFAR100-C, and TinyImageNet-C, and due to page limitations, the results for these datasets are detailed in Appendix A.4.
Figure 3: Relative robustness accuracy gain (%) on ImageNet-C (top) and ImageNet-$\bar{\mathbf{C}}$ (bottom) for ResNet50, trained using DST with gradient-based strategies (i.e., (a) RigL, (b) MESTg, (c) GraNetg) at a sparsity ratio of 0.1, compared to a dense baseline (which has a mean robustness accuracy of $38.38\%$ on ImageNet-C and $40.38\%$ on ImageNet-$\bar{\mathbf{C}}$). Positive values reflect better performance by the sparse models compared to the dense baseline. The title indicates the mean robustness accuracy for each corresponding DST method.
@@ 4.3 Robustness of DST models: is true for other datasets, tasks, and architectures?
Evaluation on video classification. We extended our exploration to video data using the UCF101 dataset Soomro et al. (2012), an action recognition dataset containing 101 categories and approximately 13K videos. Following the settings described in Schiappa et al. (2023), we compared the performance of dense training and DST at a sparsity level of $0.1$ on two architectures: 3D ResNet50 and Two-Stream Inflated 3D ConvNet (I3D) Carreira & Zisserman (2017). Both models were trained from scratch for 100 epochs on UCF101. Our analysis, based on a test set featuring 16 corruptions across severity levels 1 to 4, as shown in Table 1, revealed that DST models consistently achieved superior robust accuracy compared to their dense counterparts, even within the context of video data.
@@
In summary, our empirical findings indicate that DST methods can surpass Dense Training models in robustness performance across various datasets at specific sparsity ratios, without (or even reduce) extra resource cost  [^4] We use a binary masking operation to simulate the sparisty, the computation cost of binary masking operation is negligible or even reduced when the hardware support sparse matrix product.. Furthermore, DST models tend to perform better against different types of corruption, particularly those with more high-frequency information. The mystery behind these observations will be further explored in the following sections.
@@ 5.2 Analysis From Spectral Domain
Figure 7: (Left) A toy example demonstrating the impact of high- and low-frequency attenuation on dense ResNet50 model using the ImageNet test set. (Right-top) High-frequency attenuation performance: (a) ResNet34 (sparsity $0.5$) on CIFAR100, (b) EfficientNet-B0 (sparsity $0.5$) on TinyImageNet, and (c) ResNet50 (sparsity $0.1$) on ImageNet. (Right-bottom) Low-frequency attenuation performance for the same models and datasets. The x-axis represents the attenuation radius, and the y-axis shows test accuracy.

### v2

@@ Abstract
It is generally perceived that Dynamic Sparse Training opens the door to a new era of scalability and efficiency for artificial neural networks at, perhaps, some costs in accuracy performance for the classification task. At the same time, Dense Training is widely accepted as being the “de facto” approach to train artificial neural networks if one would like to maximize their robustness against image corruption. In this paper, we question this general practice. Consequently, we claim that, contrary to what is commonly thought, the Dynamic Sparse Training methods can consistently outperform Dense Training in terms of robustness accuracy, particularly if the efficiency aspect is not considered as a main objective (i.e., sparsity levels between 10% and up to 50%), without adding (or even reducing) resource cost. We validate our claim on two types of data, images and videos, using several traditional and modern deep learning architectures for computer vision and three widely studied Dynamic Sparse Training algorithms. Our findings reveal a new yet-unknown benefit of Dynamic Sparse Training and open new possibilities in improving deep learning robustness beyond the current state of the art.
@@ 1 Introduction
If one would like to maximize as much as possible the robustness of a deep learning model on noisy more realistic data (e.g., image corruption, out of distribution), what would be the typical approach? Would one train directly a dense neural network (referred further as Dense Training) or a sparse neural network (referred further as Sparse Training)? Of course, there are many techniques to enhance the model robustness, but still they would be applied on top of a basic optimization process which usually requires training an artificial neural network. In order to give an answer to the previously posed question, we would need to quantify the research and practitioner community perception, which is a hard problem in itself. Yet, given the large body of literature on the topic, it is safely to assume that the usual answer would be “Yes, Dense Training is the typical solution to maximize the model robustness.”. Nevertheless, the reader can answer to this question by themselves, while further in this manuscript we would like to bring more clarity and some new perspectives on this topic.
@@
DST shares many similarities with dense training, including its adaptability to a range of learning paradigms, such as continual learning Sokar et al. (2023) and reinforcement learning Graesser et al. (2022), as well as tasks like feature selection Atashgahi et al. (2022), time series classification Xiao et al. (2022). It is compatible with both convolutional and transformer-based architectures and uses core techniques such as dropout, weight decay, and standard optimizers Dey et al. (2024). Despite these parallels, DST utilizes hard sparsity—on top of the soft sparsity introduced by weight decay or other regularization methods—which provides additional regularization. This naturally leads to an intriguing question: Is somehow DST implicitly improving model robustness, and no one knew about this yet?
Although several studies have investigated the impact of pruning model from Dense Training, lottery ticket hypothesis methods Frankle & Carbin (2019); Frankle et al. (2020); Renda et al. (2020) on image corruption robustness Liebenwein et al. (2021); Diffenderfer et al. (2021) and adversarial robustness Liao et al. (2022), other research shows that static sparse networks, when increased width and depth to maintain capacity, can also improve robustness Timpl et al. (2022). However, a closer look at current DST research reveals a literature gap: the overlooked evaluation of DST models under various unexpected image corruptions. In an attempt to respond to this, we examined one of the classical DST algorithms, Sparse Evolutionary Training (SET) Mocanu et al. (2018), as illustrated in Figure 1 (Left). We find that SET (with a sparsity ratio of $0.5$) can markedly improve model robustness against 18 out of 19 image corruptions, compared to Dense Training. Intuitively, adding regularization (e.g., weight decay or dropout) during training is commonly believed to help improve the generalization of neural networks. In light of the strong performance of the DST method (e.g., SET) against common corruptions , we propose the Dynamic Sparsity Corruption Robustness (DSCR) Hypothesis.
@@ 2.2 Dynamic Sparse Training
Dynamic Sparse Training (DST) Mocanu et al. (2018); Bellec et al. (2018); Mostafa & Wang (2019); Evci et al. (2020); Chen et al. (2021); Yuan et al. (2021); Wu et al. (2025), as a sparse-to-sparse training paradigm, starts from scratch and maintains partial parameters throughout training, providing a promising approach to significantly improve training and inference efficiency while still preserving accuracy. Recent research has focused on deciphering the mechanisms of DST Liu et al. (2021c); Evci et al. (2022; 2019), and has broadened its application in various domains, such as continuous learning Sokar et al. (2023), reinforcement learning Graesser et al. (2022); Tan et al. (2023), feature selection Sokar et al. (2022); Atashgahi et al. (2022), time series classification Xiao et al. (2022), network architecture design Liu et al. (2023), and vision transformers Chen et al. (2021).
Moreover, recent studies have concentrated on developing improved DST algorithms Dettmers & Zettlemoyer (2019); Mostafa & Wang (2019); Yuan et al. (2021); Jayakumar et al. (2020); Liu et al. (2021a). For instance, Dettmers & Zettlemoyer (2019) introduced a technique to regrow weights based on the momentum of existing weights, while Yuan et al. (2021) proposed a method that utilizes both weight and gradient information for pruning. More recently, it has been demonstrated that DST can enhance the performance of DNNs on biased data Zhao et al. (2023) and adversarial samples Chen et al. (2022b). However, the majority of DST research relies on pristine benchmark datasets, leaving the robustness of DST models against image corruption relatively unexplored. This paper aims to systematically investigate the robustness behavior of DST when exposed to various image corruptions.
@@ 3.1 Problem Definition
Given $n_{\text{tr}}$ i.i.d. training samples $\left\{\left(\bm{x}_{i}^{\operatorname{tr}},y_{i}^{\operatorname{tr}}\right)\right\}_{i=1}^{n_{\text{tr }}}$ and $n_{\text{te}}$ test samples $\left\{\left(\bm{x}_{i}^{\text{te }},y_{i}^{\text{te}}\right)\right\}_{i=1}^{n_{\text{te}}}$, we aim to evaluate the model’s robustness in scenarios where the test inputs are corrupted relative to the benchmarked training datasets, (i.e., $p_{\text{tr}}(\bm{x})\neq p_{\text{te}}(\bm{x})$) while the conditional input-label relationship remains consistent (i.e., $\left.p_{\operatorname{tr}}(y\mid\bm{x})=p_{\mathrm{te}}(y\mid\bm{x}\right)$). To achieve this, we utilize standard image benchmark datasets such as CIFAR10 Krizhevsky et al. (2009), CIFAR100 Krizhevsky et al. (2009), TinyImageNet Le & Yang (2015) and ImageNet Deng et al. (2009) for training. Then, we evaluate the performance (robustness accuracy) of the models on the corresponding corruption benchmark datasets, including CIFAR10-C, CIFAR100-C, TinyImageNet-C and ImageNet- C/$\bar{\mathbf{C}}$/3DCC. More details of datasets can be found in Appendix A.2.
@@ 3.2 Dynamic Sparse Training
  • DST with soft memory budget in forward: Memory-Economic Sparse Training (MEST) Yuan et al. (2021), with Soft Memory Bound Elastic Mutation, starts with a sparse model having a memory budget of $\frac{\|\bm{\theta}^{\prime}\|_{0}}{\|\bm{\theta}\|_{0}}=b+b_{s}$, where $b_{s}>0$ represents the soft memory bound that decays during training. Subsequently, at intervals of $\Delta t$, weights are removed based on magnitude and gradient information, reducing the memory budget to $b$, and then regrown it to $b+b_{s}$, as illustrated in Figure 1 (right (c)). The soft memory constraint $b_{s}$, gradually decays to zero, allowing the model to reach the target memory budget $b$.
@@ 4.1 Overall Performance
Figure 2 showcases the average robust accuracy across various types of corruption, with severity levels ranging from 1 to 5, on the CIFAR10/100-C and TinyImageNet-C datasets. We can find that DST models outperform their Dense training counterparts (as indicated by the blue line in Figure 2) in terms of robustness accuracy across several sparsity ratios. This trend is evident in both the random regrow (Figure 2, left) and gradient-based regrow (Figure 2, right) strategies. Specifically, for CIFAR100-C and TinyImageNet- C, DST models with sparsity ratios ranging from $0.3$ to $0.7$ generally show encouraging improvements in robustness. For CIFAR10-C, at certain sparsity ratios, such as $0.4$, the overall DST methods achieve decent robust performance against image corruptions. Moreover, based on our experimental setup, at least $40\%$ of both training computational and memory costs can be saved while simultaneously enhancing corruption robustness. Details on FLOPs and parameter counts are available in the Appendix A.3.2. This finding substantiates our DSCR Hypothesis across different DST methods, varying sparsity ratios, and architectures (i.e., VGG, ResNet, and EfficientNet).
@@ 4.2 Performance on Corruption Types
We further evaluate the DST methods across various corruption types with different severities. Figure 3 provides detailed relative robustness gains  [^2] Defined as $({\tt Acc_{s,sl}-\tt Acc_{d,sl}})/{\tt Acc_{d,sl}}$, where $\tt Acc_{s,sl}$ and $\tt Acc_{d,sl}$ represent the accuracies of models trained using DST and dense training, respectively, at a severity level of $\tt sl$. for each type of corruption in ImageNet-C and ImageNet-$\bar{\mathbf{C}}$ Mintun et al. (2021), compared to their dense model counterparts. It is worth mentioning that the X-axis in Figure 3 represents different types of corruption, which are ordered according to the amount of high-frequency information they contain, as in Saikia et al. (2021). Intriguingly, on ImageNet-C and ImageNet-$\bar{\mathbf{C}}$  [^3] Our experiments primarily focus on DST with gradient-based regrow strategy when training on ImageNet. As our implementation is based on PyTorch, this regrow approach tends to achieve strong performance on ImageNet more quickly with this framework Liu et al. (2021a). However, it is worth noting that the random-based approach has also shown significant improvements when implemented in JAX Lee et al. (2024), which can be future exploration., DST models show significant potential in enhancing robustness against specific corruptions, such as impulse, Gaussian, and shot noise on ImageNet-C, as well as brownish, Perlin, plasma, and blue noise on ImageNet-$\bar{\text{C}}$. It is worth noting that these corruptions are typically characterized by relatively high-frequency information. This finding is consistent on CIFAR10-C, CIFAR100-C, and TinyImageNet-C, and due to page limitations, the results for these datasets are detailed in Appendix A.4.
Figure 3: Relative robustness accuracy gain (%) on ImageNet-C (top) and ImageNet-$\bar{\mathbf{C}}$ (bottom) for ResNet50, trained using DST with gradient-based strategies (i.e., (a) RigL, (b) MESTg, (c) GraNetg) at a sparsity ratio of 0.1, compared to a dense baseline (which has a mean robustness accuracy of $38.38\%$ on ImageNet-C and $40.38\%$ on ImageNet-$\bar{\mathbf{C}}$). Positive values reflect better performance by the sparse models compared to the dense baseline. The five bars, ranging from light to dark shades, represent corruption severity levels from 1 to 5 for each type of corruption. The title indicates the mean robustness accuracy for each corresponding DST method.
@@ 4.3 Robustness of DST models: is true for other datasets, tasks, and architectures?
Evaluation on video classification. We extended our exploration to video data using the UCF101 dataset Soomro et al. (2012), an action recognition dataset containing 101 categories and approximately 13K videos. Following the settings described in Schiappa et al. (2023), we compared the performance of dense training and DST at a sparsity level of $0.1$ on two architectures: 3D ResNet50 and Two-Stream Inflated 3D ConvNet (I3D) Carreira & Zisserman (2017). Further implementation details are provided in Appendix A.3.1. Both models were trained from scratch for 100 epochs on UCF101. Our analysis, based on a test set featuring 16 corruptions across severity levels 1 to 4, as shown in Table 1, revealed that DST models consistently achieved superior robust accuracy compared to their dense counterparts, even within the context of video data.
@@
In Appendix A.9, we compare the robustness of dense trained and DST models on an image segmentation task and find that DST models are robust not only in image and video classification but also in image segmentation. In summary, our empirical findings indicate that DST methods can surpass Dense Training models in robustness performance across various datasets at specific sparsity ratios, without (or even reduce) extra resource cost  [^4] We use a binary masking operation to simulate the sparisty, the computation cost of binary masking operation is negligible or even reduced when the hardware support sparse matrix product.. Furthermore, DST models tend to perform better against different types of corruption, particularly those with more high-frequency information. The mystery behind these observations will be further explored in the following sections.
@@ 5.2 Analysis From Spectral Domain
Figure 7: (Left) A toy example demonstrating the impact of high- and low-frequency attenuation on dense ResNet50 model using the ImageNet test set. (Right-top) High-frequency attenuation and (Right-bottom) Low-frequency attenuation performance: (a) ResNet34 (sparsity $0.5$) on CIFAR100, (b) EfficientNet-B0 (sparsity $0.5$) on TinyImageNet, and (c) ResNet50 (sparsity $0.1$) on ImageNet. The x-axis represents the attenuation radius, and the y-axis shows test accuracy.
@@
Previous works Li et al. (2021); Grabinski et al. (2022) have pointed out that high-frequency components can cause frequency aliasing when downsampling operations (e.g., AveragePooling and MaxPooling) are performed in the spatial domain of neural networks, as explained by the sampling theorem in information theory. Specifically, (Li et al., 2021) demonstrated that relying on low-frequency components of wavelets improves noise robustness, attributing this to the suppression of aliasing effects. Furthermore, (Grabinski et al., 2022; 2023) showed that removing high-frequency components to reduce frequency aliasing enhances the robustness of networks. These studies validate that reducing high-frequency components mitigates frequency aliasing and improves robustness, aligning with our findings on the relationship between DST models’ tendency to pay less attention to high-frequency information and their improved robustness against common corruptions.